# Revenue Maximization Under Sequential Price Competition Via The Estimation Of $s$-Concave Demand Functions

### v3

@@ Appendix I Remarks
where $\tau_{\text{base}}=\tfrac{2}{3}T^{5/7}$. The allocation parameter $\kappa_{i}$ for the first part of exploration is computed using Equation˜14, with $\tau$ replaced by $\tau_{i}$ and fixed constants $\mathscr{C}_{i}=\mathscr{B}_{i}=1$. For this simulation, we set $\beta_{i}=0.8$ for $N=2$, $\beta_{i}=0.92$ for $N=4$, and $\beta_{i}=97$ for $N=6$, which satisfy ˜3.6.
@@
For each horizon $T\in\{100,400,800,1600,3200\}$, we run Algorithm˜1 and repeat the simulation 30 times, reporting averages and $95\%$ confidence intervals.
The results are reported in Figure˜7, together with a log-log plot in Figure˜6, which illustrate the convergence of prices to the Nash equilibrium and the decay of regret.
As anticipated, the estimation of $\psi_{i}$ deteriorates, in line with the intuition of Remark˜4.1. Indeed, as the horizon $T$ grows, the exploration lengths $\tau_{i}$ increase for all sellers, which enlarges the non-common exploration phase (see Figure˜7). Consequently, a larger number of price vectors $\mathbf{p}^{(t)}$ are no longer i.i.d., making parameter estimation increasingly inconsistent. Nevertheless, despite the adverse effect on the estimation of $\psi_{i}$, both convergence to the Nash equilibrium and regret performance remain robust, with empirical regret rates matching – or even surpassing – the theoretical benchmarks of 1/7-1/7 for equilibrium convergence and $5/7$ for regret.

### v4

@@ Appendix I Remarks
where $\tau_{\text{base}}=\tfrac{2}{3}T^{5/7}$. The allocation parameter $\kappa_{i}$ for the first part of exploration is computed using Equation 14, with $\tau$ replaced by $\tau_{i}$ and fixed constants $\mathscr{C}_{i}=\mathscr{B}_{i}=1$. For this simulation, we set $\beta_{i}=0.8$ for $N=2$, $\beta_{i}=0.92$ for $N=4$, and $\beta_{i}=97$ for $N=6$, which satisfy 3.6.
@@
For each horizon $T\in\{100,400,800,1600,3200\}$, we run Algorithm 1 and repeat the simulation 30 times, reporting averages and $95\%$ confidence intervals.
The results are reported in Figure 7, together with a log-log plot in Figure 6, which illustrate the convergence of prices to the Nash equilibrium and the decay of regret.
As anticipated, the estimation of $\psi_{i}$ deteriorates, in line with the intuition of Remark 4.1. Indeed, as the horizon $T$ grows, the exploration lengths $\tau_{i}$ increase for all sellers, which enlarges the non-common exploration phase (see Figure 7). Consequently, a larger number of price vectors $\mathbf{p}^{(t)}$ are no longer i.i.d., making parameter estimation increasingly inconsistent. Nevertheless, despite the adverse effect on the estimation of $\psi_{i}$, both convergence to the Nash equilibrium and regret performance remain robust, with empirical regret rates matching – or even surpassing – the theoretical benchmarks of 1/7-1/7 for equilibrium convergence and $5/7$ for regret.
@@
J.3 Robustness to misspecification of $s_{i}$.
In this experiment, we replicate the setup of Section J.1, except that we fix $N=4$ and set $\beta_{i}=0.9$. We again take $\psi_{i}(u)=\Phi(2u/i)$ for $i=1,2,\dots,N$, where $\Phi$ denotes the c.d.f. of a standard Gaussian random variable. Since $\Phi$ is log-concave, each $\psi_{i}$ is 0-concave, that is, $s_{i}=0$. We then run Algorithm 1 under five different shape parameters $s_{i}^{\prime}\in\{-0.2,-0.1,0,0.1,0.2\}$. As shown in Figure 8, this misspecification of $s_{i}$ has no visible effect on either the NE convergence rate or the regret convergence rate. The small variations in the estimated slopes reported in the legend of Figure 8 are plausibly due to finite-sample noise ($T=1600$).
Figure 8: Log-log performance of Algorithm 1 in sequential price competition with $N=4$ sellers for different values of misspecification of $s_{i}=0$, specifically $\{-0.2,-0.1,0,0.1,0.2\}$. The slopes, indicated as $m$, of the convergence to NE and the regret are close to each other. The small variations in the rates can be attributed to the finite sample experiment ($T=1600$).